### v3

# Flat bands and unconventional superconductivity in a simple model of metal-organic frameworks

Abstract
We show that interference-induced flat bands (IFBs) are commonplace in reticular materials. For example, several metal-organic frameworks, including the superconductor Cu-BHT, form kagome lattices with metals at the vertices and ligands along the bonds. Similar bipartite motifs are common in reticular materials. A tight-binding model on this lattice yields partially occupied IFBs at half-filling with large gaps between them and all other bands. Long-range hopping induces curvature in the bands but leaves them flatter and more isolated than those in twisted bilayer graphene. The slave boson theory of the $t$-$J$ model on this lattice shows unconventional superconductivity. Thus, crystal engineering provides a lattice-driven route to flat bands with high electronic densities.
Reticular materials are crystalline frameworks built by linking molecular building blocks with strong (e.g., covalent, dative) bonds; metal-organic frameworks (MOFs), coordination polymers (CPs) and covalent organic frameworks (COFs) provide prominent examples [1]. Many correlated insulating states occur in reticular materials [2, 3, 4, 5, 6, 7, 8, 9, 10]. Therefore, it has been argued that these systems may display unconventional superconductivity when driven out of their insulating phases, e.g., by doping [11], pressure, or control of chemistry [12]. This hypothesis received strong support from the recent discovery of superconductivity in copper(II) benzenehexathiolate (Cu-BHT) [13], and from the evidence of unconventional superconductivity and strong electronic correlations in this material [14].
Contemporaneously, the discovery of superconductivity and strongly correlated insulating states in twisted bilayer graphene (TBG) revived interest in flat band superconductivity [15, 16]. In TBG, a group of isolated ‘nearly flat bands’, with a total bandwidth on the order of $10$ meV, and band gaps of the same order on either side [15, 17], cross the Fermi energy. In a flat band, the kinetic energy of all electrons is equal (hence, the density of states is enormous). Thus, interactions determine the behavior of the system, amplifying strongly correlated phenomena, such as the fractional quantum Hall effect, high temperature superconductivity, charge density waves, and metal-insulator transitions [18, 19, 20, 21, 22]. However, TBG has extremely low electron densities, and many of the most interesting predictions for flat bands require high electron densities [23, 18]. Thus, discovering bulk materials with flat bands is an important goal.
Reticular materials offer the unique possibility of “crystal engineering” [24, 25]: realizing structures by design. Therefore, an important task for theory is to identify structures that promise the realization of new physical phenomena. We show below that reticular materials are ideally suited for the study of flat band physics.
In this Letter, we argue that reticular materials generically form bipartite lattices with flat bands, which can be understood as a consequence of Sutherland’s theorem [26, 27]. We explore a simple example; the kagome-Lieb lattice, Fig. 1a, which is to the kagome lattice as the Lieb lattice is to the square lattice. We show that an $s$-orbital model on this lattice has five interference-induced flat bands (IFBs), three of which are partially occupied at half-filling, with large gaps between them and all other bands Fig. 1b. We show that, and explain why, they are robust to the largest expected perturbations, Figs. 2, 3. The multi-orbital analog has a closely related band structure with IFBs, Fig. 4, which can be understood in terms of the s-orbital model, and reproduces the key features of the band structure of the superconducting MOF, Cu-BHT, in which the metals form a kagome lattice with ligand sites along the bonds. When strong electronic correlations are included, we find that the $s$-orbital model leads to three degenerate superconducting phases, Fig. 5.
Figure 1: Tight-binding model on the kagome-Lieb lattice. (a) Hopping integrals are marked, $\delta$ is the site energy of the metals, and the shaded area is the unit cell. The Bravais lattice is inset; $\Gamma$, $M$ and $K$ label the high symmetry points. (b) Band structure of the nearest-neighbor model.
CPs consist of metals bonded to organic groups (ligands) [1]. These ligands link to other metals to form a crystal. The large unit cells of CPs mean that beyond nearest-neighbor hopping is often negligible, resulting in bipartite lattices.
For a $d$-dimensional CP with periodic boundary conditions, $|N_{L}-N_{M}|\geq d-1$, where $N_{L}$ ($N_{M}$) is the number of ligand (metal) sites per unit cell [28]. Whence Sutherland’s theorem [26] requires that the nearest-neighbor tight-binding model for a reticular material must contain at least $d-1$ degenerate flat bands. This makes reticular materials playgrounds for flat bands.
MOFs are the subclass of CPs with high porosity [29], and hence have widespread potential applications, such as carbon capture and storage, hydrogen fuel storage, microelectronics, and drug delivery [29, 2, 3]. COFs are similar to CPs and MOFs, but with the metals replaced by an organic group [30, 4]. Therefore, one also expects zero energy flat bands in MOFs and COFs. ‘Crystal engineering’ [24, 25]: the rational design of the crystal structures of reticular materials is enabled by the local nature of chemical bonding [1, 31, 24, 25, 32]. Thus, realizing specific lattice structures in reticular materials is a highly achievable objective. Reticular lattices could also be realized in optical lattices.
The kagome-Lieb lattice.-To substantiate these ideas, we explore an example. In many CPs and MOFs [13, 14, 33, 34, 35, 36, 37, 38] the metals form a kagome lattice with the ligands lying along the nearest-neighbor bonds. For example, density functional theory (DFT) calculations show that the low-energy electronic structure of Cu-BHT and is dominated by bands that originate from the Cu and S atoms with little weight on the C atoms [12]; structurally, the S atoms sit along the nearest-neighbor Cu-Cu bonds. As both the metals and ligands have orbitals near the Fermi energy and the s-orbital tight-binding approximation is applicable to many MOFs [39, 40, 41, 42], the simplest possible model is the tight-binding model on the kagome-Lieb lattice, Fig. 1a. Importantly, increasing covalency between metals and ligands is a design strategy for achieving metallic and superconducting MOFs [2], which motivated the first studies of Cu-BHT [14], and necessitates electronic models including both the metals and the ligands. This lattice also occurs quite naturally in COFs [30]. In contrast, the kagome-Lieb lattice is absent from recent catalogs of flat-band inorganic materials [23, 43, 44]. Thus, we study the Hamiltonian
$H_{tb}=-\sum_{i,j,\mu,\nu,\sigma}t_{ij}^{\mu\nu}\hat{c}_{i\mu\sigma}^{\dagger}\hat{c}_{j\nu\sigma}+\sum_{i\in M,\mu,\sigma}\delta_{\mu}\hat{c}_{i\mu\sigma}^{\dagger}\hat{c}_{i\mu\sigma},$ (1)
where $\hat{c}_{i\mu\sigma}^{(\dagger)}$ creates (annihilates) an electron of spin $\sigma$ in orbital $\mu$ at site $i$, $t_{ij}^{\mu\nu}$ is the hopping amplitude between orbitals on different sites, and the sum over $i\in M$ includes metal sites only.
Generically, reticular materials will be multi-orbital, but we can learn a lot about the kagome-Lieb lattice by considering the simplest model with one $s$-orbital per site: we suppress the orbital indices and sketch the hopping amplitudes $t_{ij}^{\mu\nu}\in\{t,t_{M},t_{LH},t_{LT}\}$ in Fig. 1a.
In the nearest-neighbor model ($\delta=t_{M}=t_{LH}=t_{LT}=0$) there are five IFBs: three degenerate IFBs at $\varepsilon_{\bm{k}}=0$ and two ‘kagome-like’ IFBs, which touch kagome-like dispersive bands (Fig. 1b). The three degenerate flat bands are required by Sutherland’s theorem [26] for $N_{L}=6$, $N_{M}=3$; however, the kagome-like flat bands are not.
Figure 2: (Unnormalized) localized eigenstates for each of the interference-induced flat bands. Electrons cannot hop outside the loops due to destructive interference. The (a) antibonding ($|+\rangle_{h}$) and (b) bonding ($|-\rangle_{h}$) kagome-like eigenstates have energies $E_{\pm}=[t_{\text{LT}}+2t_{\text{M}}+\delta\pm\sqrt{8t^{2}+(2t_{\text{M}}-t_{\text{LT}}+\delta)^{2}}]/2$ and a relative amplitude on the metal sites of $\alpha_{\pm}=\left(E_{\pm}-t_{\text{LT}}\right)/t$ (for $t_{\text{LH}}=0$). The three non-bonding (between metals and ligands) eigenstates (c-e) can be constructed as superpositions of bow-tie states (f), as can noncontactable loops for PBC (g). The superpositions are indicated by the arrows which define the sign of the bow-tie wavefunctions centered on each metal site.
Localised states.-The five flat bands result from quantum interference due to the topology of the lattice and are thus IFBs [19, 45, 46, 47, 48]. The localized states that comprise the flat-band of the kagome lattice consist of alternating sign amplitudes around each hexagon [45]. The kagome-like IFBs on the kagome-Lieb lattice are composed of the (anti)bonding combination of the metals and two nearest-neighbor ligands on hexagon $h$. These hybrid orbitals form flat bands, ($|+\rangle_{h}$) $|-\rangle_{h}$, analogous to those found in the kagome lattice (Fig. 2a,b). The dispersive bands with (high) low energies are also kagome bands in the basis of the (anti)bonding states of metals/nearest-neighbor ligands. This explains the Dirac point at K (Fig. 1b). In contrast, the three degenerate IFBs are non-bonding (between metals and ligands), as they have no weight on the metal sites (Fig. 2c-e). The non-bonding states can also be understood as superpositions of ‘bow-tie’ states Fig. 2f.
Figure 3: Evolution of the band structure of the kagome-Lieb lattice tight-binding model with hopping beyond nearest-neighbors. Five bands remain exactly flat even when (a) metal-ligand orbital energy differences, $\delta$, (b) direct metal-metal hopping, $t_{\text{M}}$, or (c) both are included. (d) Ligand-ligand hopping across triangles, $t_{\text{LT}}$, leaves three exactly flat bands. (e) Ligand-ligand hopping across hexagons, $t_{\text{LH}}$, leaves one flat band. (f) Including all of these terms leaves no bands exactly flat, but the non-bonding bands remain gapped and narrow. The bandwidth of the non-bonding bands, $W$, and the electron (hole) bandgap, $E_{g}^{(e)}$ ($E_{g}^{(h)}$), are labeled. (g,h) Even for large $t_{M},t_{LT}$ and $t_{LH}$, the non-bonding bands remain much flatter and more isolated (i.e., have larger $E_{g}/W$) than the nearly-flat bands in TBG [17].
The two kagome-like flat bands are singular as they touch a dispersive band at the $\Gamma$ point ($k=0$). This has a simple explanation, analogous to that for the kagome [45] and other decorated kagome lattices [47]. Each unit cell contains one hexagon (Fig. 1). Therefore, for $N$ unit cells and open boundary conditions (OBC) there are $N$ (anti)bonding kagome-like IFB states. Superpositions of |±$|\pm\rangle_{h}$ states on adjacent hexagons have zero weight on the sites that contribute to both localized wavefunctions due to destructive interference. Therefore, $\sum_{h}|+\rangle_{h}=\sum_{h}|-\rangle_{h}=0$ for periodic boundary conditions (PBC), leaving $N-1$ non-trivial superpositions of $|+\rangle_{h}$ ($|-\rangle_{h}$). However, for PBC one can construct two additional states, which cannot be written as superpositions of $|+\rangle_{h}$ ($|-\rangle_{h}$), and wrap around the lattice (i.e., one noncontractable loop composed of (anti)bonding states in each of the $x$ and $y$ directions) [45, 47]; resulting in $N+1$ states with energy $E_{\pm}$. As each IFB can only accommodate $N$ states, topology dictates that there cannot be a gap between the kagome-like IFBs and the dispersive bands.
The non-bonding states are most easily counted in the basis of bow-tie states (Fig. 2f). There are three metal sites, and thus three linearly independent bow-tie states, per unit cell. All noncontractable loops are superpositions of bow-ties (Fig. 2g). Thus, the $3N$ bow-tie states are non-singular and need not touch a dispersive band.
The non-singularity of the non-bonding IFBs allows significant gaps between the non-bonding IFBs and the dispersive bands. In contrast, the IFBs on the kagome [45], Lieb [46], pyrochlore [45], and star [47] lattices are all singular and have topologically required band touchings, and therefore, excitations into dispersive bands occur at arbitrarily low energies. Thus, materials realizing these lattices are significantly less promising candidates to explore flat band physics at high electronic densities than materials that form kagome-Lieb lattices. New reticular materials tailored to exploit this structure could be readily synthesized to complement existing examples, such as Cu-BHT.
Beyond nearest-neighbor hopping.- A mismatch in the orbital energies of the metals and ligands, $\delta\neq 0$, does not induce a curvature in the degenerate IFBs as the lattice remains bipartite, (and, more intuitively, none of the localized wavefunctions contain any amplitude on the metals, Fig. 2c-e) [26]. The two kagome-like IFBs remain, as $\delta$ simply changes the amplitudes on the metal and ligand orbitals in the (anti)bonding orbitals. However, $\delta\neq 0$ breaks particle-hole symmetry; positive $\delta$ increases (decreases) the energy gap between the non-bonding IFBs and the (anti)bonding kagome-like bands, $E_{g}^{(h)}$ ($E_{g}^{(e)}$), Fig. 3a.
Direct metal-metal hopping, $t_{\text{M}}$, does not cause any of the five IFBs of the kagome-Lieb lattice to become dispersive, Fig. 3b. This is surprising as the lattice is no longer bipartite for $t_{M}\neq 0$. Therefore, Sutherland’s theorem no longer guarantees that the model retains the three non-bonding IFBs. However, as the non-bonding IFBs have zero amplitude on all metal sites, metal-metal hopping cannot perturb them. The two kagome-like IFBs have non-zero amplitudes on the metal sites. However, in $|+\rangle_{h}$ and $|-\rangle_{h}$, the phases of the metal sites alternate around each hexagon and are therefore localized by destructive interference under direct metal-metal hopping, Fig. 2. Therefore, the contribution to the energy of kagome-like IFBs from metal-metal hopping is independent of momentum and these bands remain completely flat.
Thus, the five IFBs remain completely dispersionless, regardless of the values of $\delta$ and $t_{\text{M}}$, Fig. 3a-c. This is important because the values for $\delta$ and $t_{\text{M}}$ in CPs and MOFs can vary dramatically, even between closely related materials [49].
Direct ligand-ligand hopping is typically small; however, it is not always negligible [50, 6]. We consider ligand-ligand hoppings across triangles, $t_{LT}$, and across hexagons, $t_{LH}$; Fig. 1a. For example, the latter could be mediated by virtual hoping via the S-$p_{z}$ orbitals in Cu-BHT. Both have important qualitative effects on the band structure of the kagome-Lieb lattice, Fig. 3d-f. If exactly one of $t_{LT}$ and $t_{LH}$ is non-zero only one of the non-bonding IFBs remains exactly flat. This can be understood from interference and lattice topology as a simple extension of the above arguments [28].
For both $t_{LT}\neq 0$ and $t_{LH}\neq 0$, no completely flat bands remain, Fig. 3f. However, the non-bonding bands remain nearly flat and widely gapped, indeed, significantly flatter and more isolated than the nearly flat bands in TBG, Fig. 3 g,h. This model reproduces the key features of the electronic structure of the MOF Cu3C6O6 [51], which is chemically analogous to Cu-BHT with O substituted for S.
The large unit cells of reticular materials mean that longer range hopping is likely to be negligible. Thus we expect that these conclusions will hold in many framework materials described by kagome-Lieb lattice tight-binding models.
It is also interesting to anticipate future experiments where a 2D kagome-Lieb material is grown on a substrate. Strain caused by the substrate will perturb the ideal lattice and hence the hopping paramaters. The non-bonding IFBs will be robust to such distortions because they do not change the bipartite nature of the lattice so Sutherland’s theorem still holds.
Multi-orbital tight-binding model- The multi-orbital tight-binding model (Eq. 1) with a Slater-Koster parameterization [28] reproduces the key features of the DFT band structure of Cu-BHT [12], Fig. 4. The multi-orbital tight-binding model hosts several flat and nearly flat bands, which can readily be understood from simple generalizations of the arguments above. The model is symmetric under reflection through the plane and hence is separable into even (odd) sectors, with support from ligand $\pi_{x}$ and $\pi_{y}$ orbitals, and metal $d_{xy}$, $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ (ligand $\pi_{z}$ and metal $d_{xz}$ and $d_{yz}$) orbitals. These sectors are decoupled at the single-electron level, but there are Coulomb interactions between electrons in the different sectors. The $d_{z^{2}}$ orbitals are energetically separated and do not hybridize significantly with the other orbitals in Cu-BHT; therefore we neglect them below. This leaves 12 (6) ligand and 6 (6) metal orbitals per unit cell in the even (odd) sector. Therefore, we expect 6 even non-bonding IFBs with support on the $\pi_{x}$ and $\pi_{y}$ orbitals (Fig. 4).
There are twice as many IFBs in the even sector of the multi-orbital model as there are in the $s$-orbital model because there are two orbitals per site in the even sector. The compact localised states that give rise to the IFBs in the multi-orbital model are simply related to those that give rise to the non-bonding IFBs in the s-orbital model: the single orbitals are replaced by a superposition of the two orbitals on a given site. This leads to a deep connection between the resulting IFBs in the two models: both are required by Sutherland’s theorem, are robust to strain, remain exactly flat when metal-metal hopping is added even though the lattice is no longer bipartite, and remain very narrow when ligand-ligand hopping is added [28].
Figure 4: Interference-induced nearly flat bands remain in a multi-orbital kagome-Lieb lattice model, with $d$-orbitals on the metals and $\pi$-orbitals on the ligands [28]. Here we show the model for parameters chosen to reproduce the key features of Cu-BHT (cf. Ref. [12]). $V_{pd\pi}\sim 1$ eV is a Slater-Koster energy integral [28, 52, 53].
Figure 5: (a) Three degenerate superconducting phases emerge on doping from half-filling ($s$, $s+d$ and $d+id$). Here $J=0.2t$, the inverse temperature is $\beta t=2000$, $\tilde{\Delta}$ is the superconducting order parameter, and $x$ is the hole doping per unit cell. (b) Sketches of the non-trivial superconducting order parameters.
Superconductivity.- One expects metals with partially occupied flat bands to be unstable. Various symmetry breaking or topological orders can occur, including the fractional quantum Hall effect, magnetic states, charge density waves, or superconductivity. Determining, theoretically, which is realized is extremely challenging. Nevertheless, to illustrate the role of electronic correlations, we study the nearest-neighbor $s$-orbital kagome-Lieb $t$-$J$ model:
${H_{tJ}=\hat{P}\Big{[}H_{tb}+J\sum_{\langle i,j\rangle_{1}}\Big{(}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}-\frac{1}{4}\hat{n}_{i}\hat{n}_{j}\Big{)}\Big{]}\hat{P},}$ (2)
where $\hat{P}$ is the Gutzwiller projector and $\hat{\bm{S}}_{i}$ ($\hat{n}_{i}$) is the spin (number) operator on site $i$. We write the electronic operators $\hat{c}_{i\sigma}=\hat{b}_{i}^{\dagger}\hat{f}_{i\sigma}$ in terms of (slave) bosonic holons, $\hat{b}_{i}$, and fermionic quasiparticles, $\hat{f}_{i\sigma}$ [54] and calculate the superconducting order parameter, $\tilde{\Delta}_{ij}=\frac{J}{2}\langle\hat{c}_{i\uparrow}\hat{c}_{j\downarrow}-\hat{c}_{i\downarrow}\hat{c}_{j\uparrow}\rangle$, in the mean-field approximation [55, 28].
For non-zero $\delta$, a charge transfer insulator [56] arises near half filling, due to the energy benefit of exclusively doping holes onto either the ligands or the metals, depending on the sign of $\delta$. One sublattice remains exactly half-filled, thus charge transport is forbidden by the Gutzwiller projector in Eq. 2.
Upon further hole doping we find three degenerate superconducting states: $s$ ($A_{1}$), $s+d$ ($A_{1}+E_{2}$) and $d+id$ ($E_{2}$) characters, Fig. 5b. High degeneracy is not unexpected in a flat band system, but it is interesting that it survives in the strongly correlated limit considered here. The $d+id$ breaks time reversal symmetry. The $s$ state is fully gapped. The $d+id$ and $s+d$ states have line (surface) nodes in 2D (3D) [57, 58] at optimal doping. However, these nodes are formally accidental, and lift away from optimal doping. Interestingly, there is experimental evidence for nodes in Cu-BHT [14]. Non-zero $\delta$ suppresses superconductivity at low doping, but not near optimal doping, suggesting that superconductivity in reticular materials can be robust to energy differences between the metals and ligands.
While $\tilde{\Delta}\neq 0$ indicates condensing Cooper pairs, a non-zero superfluid stiffness is required to exhibit the Meissnner effect. For an isolated set of partially filled flat bands the superfluid stiffness in the weak coupling limit is given by $D_{s}\propto\text{tr}(g)|\tilde{\Delta}|$, where $g$ is the minimal Fubini-Study quantum metric [59, 60, 61, 62, 19, 63]. The equivalent condition for strong coupling superconductivity remains an open question. We find $\text{tr}(g)=0.14$ for the parameters in Fig. 3f. Thus, we expect a non-zero superfluid stiffness in the superconducting phases. For $t=1$ eV, which is comparable to Cu-BHT (Fig. 4), and $J=0.2t$, we find $|\tilde{\Delta}|=7$ K at optimal doping, somewhat larger than the critical temperature of Cu-BHT (0.25–0.3 K) [14].
Conclusions.-This demonstrates that the very narrow, isolated bands in materials with the kagome-Lieb lattice exhibit exotic, strongly correlated phenomena, including unconventional superconductivity. Thus, the kagome-Lieb lattice is an unusually rich playground for unconventional superconductivity. This provides strong support for our central hypothesis that reticular materials are an ideal platform for investigating flat band physics. A key design principle, derived from the results above, is to choose ligands with low direct ligand-ligand overlap as this minimizes the bandwidth of the non-bonding bands.
This work was supported by the Australian Research Council (DP180101483) and MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant Number JPMXS0118069605.